OpenStudy (anonymous):

n =1 The sum of an arithmetic series with 15 terms is 555. If the last term is 9, find the first term.

OpenStudy (anonymous):

sum of AS = (1st term + last term)(total number of term)/2 555=(x+9)(15)/2 x=65

OpenStudy (anonymous):

u can also apply this method: average of terms in AP= (1st term+last term)/2 here, average= 555/15=37 last term=9, so 1st term =65